Problem 1: You run a construction firm. You have just won a contract to build a government office building Building it will require an investment of $10 million today and $5 million in one year. The government will pay you $20 million in one year upon the buildings completion. Suppose the cash flows and their times of payment are certain, and the risk-free interest rate is 10%. a. What is the NPV of this opportunity? b. How can your firm turn this NPV into cash today?

Problem 1: 


You run a construction firm. You have just won a contract to build a government office building Building it will require an investment of $10 million today and $5 million in one year. The government will pay you $20 million in one year upon the building's completion. Suppose the cash flows and their times of payment are certain, and the risk-free interest rate is 10%. 


a. What is the NPV of this opportunity? 

b. How can your firm turn this NPV into cash today?
Answer #1
Answer #2

A)

step 1) 20/ (1+0.10)= 18.1818181818 or 18.18

step 2) 10+ 5/(1+0.10)= 14.5454545455 or 14.55

step 3) 18.18- 14.55= 3.63 NPV


B) firm can borrow 18.18 today pay it back 10% interest using 20 million recieve from gov't. the firm can use 10 million of the 18.18 million to cover its costs today and save 4.55 (5/ 1.10) in the bank to earn 10% interest to cover its cost of 5.00 million (4.55* 1.10). this leaves 18.18- 10.00- 4.55= 3.63 million for the firm today.

answered by: Andrew San Andres